OpenStudy (anonymous):

How do I start this? x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21

hartnn (hartnn):

subtract 1st from 2nd Equation, you will get one equation having y and z only--->(1) subtract 3rd from 2nd Equation, you will get other equation having y and z only--->(2) then u can solve (1)and(2) to get y and z

OpenStudy (anonymous):

I'm subtracting x from the second equation?

hartnn (hartnn):

u are subtracting equations,not x like this: x + 3y - z = 2 - x - 2y + 3z = 7 ___________________ 0 + 5y - 4z = -5 got this ?

OpenStudy (anonymous):

oh! and now I take that answer and subtract it from the third equation?

hartnn (hartnn):

ummmm, nopes, you subtract 3rd equation from 2nd equation similarly and again x gets eliminated

OpenStudy (anonymous):

ok I got 0 + 0 - 2z = 14

hartnn (hartnn):

y doesn't get cancelled...... x - 2y + 3z = 7 - x + 2y - 5z = -21 ------------------- 0 -4y +8z = 28 did u understand these 2 subtractions? these are term by term subtraction.

OpenStudy (anonymous):

so all the positive signs get turned to negative ones?

hartnn (hartnn):

yes,and all negative turns into positive (x - 2y + 3z = 7) - ( x + 2y - 5z = -21)

OpenStudy (anonymous):

okay, now from here what do i do?

hartnn (hartnn):

u got 2 equations in y and z only! can u solve it ?

OpenStudy (anonymous):

y +4z = 23?

hartnn (hartnn):

-4y +8z = 28 5y - 4z = -5 can u solve these 2 ?

OpenStudy (anonymous):

do i substitute?

hartnn (hartnn):

there are several methods, which method are u comfortable with? substitution , elimination ...

OpenStudy (anonymous):

i get lost once i get to 8z = 4y + 28

hartnn (hartnn):

divide by 2 on both sides, what u get ?

OpenStudy (anonymous):

where does the 2 come from?

OpenStudy (anonymous):

4z = 2y + 14 ?

hartnn (hartnn):

8z = 4y + 28 but there is -4z in other equation, right ? so how would i make 4z from 8 z ?

OpenStudy (anonymous):

oh i see now. so now i have 4z = 2y + 14 and 5y - 4z = -5

hartnn (hartnn):

yes. so put the expression of 4z u got in 5y - 4z = -5 and find y

OpenStudy (anonymous):

y = -1?

hartnn (hartnn):

5y-2y-14=-5 3y = 14-5 3y=9 y=?

OpenStudy (anonymous):

3

hartnn (hartnn):

did u get how ?

OpenStudy (anonymous):

yeah I think so, let me try to solve the rest of it

OpenStudy (anonymous):

is it x = -2,y = 3, z = 5?

hartnn (hartnn):

absolutely correct! good work :)